OpenStudy (anonymous):

Hardly a Pre-Calc question, more Alegebra 2. What are the zeros of the function x^6-x^4-16x^2+16. Pretty sure I have the answer but can you help me through the steps of factoring by grouping?

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (anonymous):

I think the zeros are (x+1)(x-1) (x+2) (x-2) (x+i) (x-i)

jimthompson5910 (jim_thompson5910):

factor by grouping to get x^6-x^4-16x^2+16 (x^6-x^4)+(-16x^2+16) x^4(x^2-1)-16(x^2-1) (x^4-16)(x^2-1) (x^2-4)(x^2+4)(x^2-1) (x-2)(x+2)(x^2+4)(x^2-1)

jimthompson5910 (jim_thompson5910):

oh sry, factor x^2 - 1 to get (x-1)(x+1), so the whole thing factors to (x-2)(x+2)(x^2+4)(x-1)(x+1)

OpenStudy (anonymous):

So there are no imaginary zeros?

jimthompson5910 (jim_thompson5910):

yes there are

OpenStudy (anonymous):

I dont understand, what the imaginary zeros are, was i right the first time?

jimthompson5910 (jim_thompson5910):

you were close

jimthompson5910 (jim_thompson5910):

solve x^2+4 = 0

OpenStudy (anonymous):

x^2=-4

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

confused with how i square root the -4

jimthompson5910 (jim_thompson5910):

break it up like this \[\Large \sqrt{-4} = \sqrt{-1*4}\] \[\Large \sqrt{-4} = \sqrt{-1}*\sqrt{4}\]

OpenStudy (anonymous):

ahh ok, so x=+-2i

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

to get the other real roots, set them equal to 0 and solve for x

OpenStudy (anonymous):

I understand now! Thank you very much for your help! :D

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Going to close it, farewell friend.

OpenStudy (anonymous):

btw might I add that you can also solve problem easily by using synthetic division